OpenStudy (anonymous):

What is the value of the missing exponent? (5x^3)^?=125x^9

ganeshie8 (ganeshie8):

hint - 5^3 = 125

OpenStudy (anonymous):

So three!!!

ganeshie8 (ganeshie8):

you got it ! Left Hand Side : put ? =3 (5x^3)^3 5^3(x^3)^3 125 x^(3*3) 125x^9 = Right Hand Side so 3 = \(\checkmark \)

OpenStudy (anonymous):

thank you!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

ganeshie8 (ganeshie8):

yw :)